yourgrandpa:

members of a band estimated that 200 people were at their show at night. their manager told them that 227 people had actually been at the show. what was the band's percent error? 200 should be the denominator. Not 227

pooja195:

Any ideas on where to start?

yourgrandpa:

nope

pooja195:

To calculate percent error you would do the following: Estimate-Actual/Actual

pooja195:

Actual= amount of people that actually came

yourgrandpa:

227 actually came

pooja195:

Correct what was the estimated?

yourgrandpa:

200

pooja195:

Good so, 200-227=?

AZ:

`200 should be the denominator. Not 227` Is this something that your teacher told you? I'm just trying to understand if you got the question wrong because this was the exact question we answered an hour ago.

yourgrandpa:

−27

yourgrandpa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ `200 should be the denominator. Not 227` Is this something that your teacher told you? I'm just trying to understand if you got the question wrong because this was the exact question we answered an hour ago. \(\color{#0cbb34}{\text{End of Quote}}\) yeah she said it was wrong

animeweeb:

mabye all the seats where filled so they didnt know there where more people

pooja195:

Is there information we are missing then because this is how you would find percent error

yourgrandpa:

this is what she said 200 should be the denominator. Not 227 that's what the teacher said

pooja195:

Can you ask for further clarification so we can better understand what she is trying to say? Because: https://www.mathsisfun.com/numbers/percentage-error.html

AZ:

Okay so \(\text{% error} = \dfrac{|\text{actual value - expected value}|}{\text{expected value}} \times 100\)

yourgrandpa:

ok

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 Can you ask for further clarification so we can better understand what she is trying to say? Because: https://www.mathsisfun.com/numbers/percentage-error.html \(\color{#0cbb34}{\text{End of Quote}}\) Yeah, I don't know. I've been seeing formulas that go both ways now.

yourgrandpa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ \(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 Can you ask for further clarification so we can better understand what she is trying to say? Because: https://www.mathsisfun.com/numbers/percentage-error.html \(\color{#0cbb34}{\text{End of Quote}}\) Yeah, I don't know. I've been seeing formulas that go both ways now. \(\color{#0cbb34}{\text{End of Quote}}\) i will ask

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Okay so \(\text{% error} = \dfrac{|\text{actual value - expected value}|}{\text{expected value}} \times 100\) \(\color{#0cbb34}{\text{End of Quote}}\) Like you've already said: `227 actually came` 200 were estimated/expected \(\text{% error} = \dfrac{\text{227 - 200}}{\text{200}} \times 100\text{%}\)

yourgrandpa:

−2700 ?

AZ:

I'm not sure how you got that Do it one step at a time 227-200 = ? Then divide by 200 Then multiply by 100

yourgrandpa:

13.500 ??

AZ:

And that's the percent error! 13.5%

yourgrandpa:

thanks once again (:

AZ:

It was my pleasure like always (: